# image generation; decoupled cross-attention; latent diffusion model; controllability

### v3

@@ 4.4 Visual Assessment
Figure 6: Sample questions for a user survey. Each question presents four options, one of which is an image generated by this study. The questions evaSluate the images’ overall aesthetic appeal, style consistency, creativity, and detail quality.
@@ 5 Discussion
6 Conclusions
Traditional hand-drawing methods for creating landscape paintings not only require the artist to have extensive experience and creativity but also involve labor-intensive efforts. This often leads to a lack of creativity and low design efficiency. To address these issues, we propose LPGen for the automatic generation of landscape paintings. Experimental results demonstrate that LPGen is a high-fidelity, controllable model capable of replacing the tedious creative and drawing tasks in the traditional design process. This transformation of the conventional landscape painting creation workflow significantly enhances design efficiency and creativity.
Nevertheless, this study acknowledges certain limitations that warrant further investigation. Firstly, controlling the generation of landscape paintings using canny edge maps and reference style images is challenging. Future improvements could be made by incorporating additional structural control methods such as depth maps, sketches, and recoloring. Secondly, while generating styles, the current method controls the style of the entire generated image, but it cannot specify the texture appearance in specific regions of the image. Therefore, enhancing the controllability of the generation process is necessary. Additionally, our understanding of painting styles is manually classified; a more extensive and automated annotation classification method might better represent real-world style classifications. Finally, the style metrics of the images generated by the proposed method can still be improved to enhance the realism and usability of the generated images.
AI can be applied to various stages of future painting creation. For example, in the conceptualization stage, AI can provide artists with a range of creative inspirations and style references to choose from. Subsequently, image generation technologies based on GANs or DMs can quickly produce numerous sketches and initial compositions. In the detail refinement stage, automated AI algorithms can optimize lines, colors, and textures, thereby reducing the workload of traditional manual retouching. Finally, automated style transfer and high-resolution rendering technologies can enhance the efficiency from creation to presentation, ultimately forming a more comprehensive AI-assisted painting creation workflow.
[…]
Conceptualization, W.Y. and Y.Z.; methodology, W.Y.; validation, W.Y., X.W. and Y.Q.; formal analysis, X.W. and Y.Q.; investigation, X.W. and Y.Q.; data curation, X.W. and Y.Q.; writing—original draft preparation, W.Y.; writing—review and editing, Y.Z.; supervision, Y.Z.; project administration, Y.Z.; funding acquisition, W.Y. All authors have read and agreed to the published version of the manuscript.
[…]
This work is jointly supported by the Beijing Municipal High-Level Faculty Development Support Program (Grant No. BPHR202203072).
[…]
The dataset used in this study will be made publicly available at a later stage and can be requested by contacting ywg@bigc.edu.cn.
Acknowledgements.
[…]

### v4

@@ IV-D Visual Assessment
Figure 6: Sample questions for a user survey. Each question presents four options, one of which is an image generated by this study. The questions evaluate the images’ overall aesthetic appeal, style consistency, creativity, and detail quality.
@@ V Discussion
VI Conclusions
This paper presents LPGen, a novel diffusion-based model that addresses the challenge of generating high-fidelity landscape paintings with a balanced control over structure and style. By introducing a decoupled cross-attention mechanism, LPGen effectively processes structural and stylistic features independently, reflecting the layered techniques used in traditional painting. The integration of a structural controller further enhances the model’s ability to maintain aesthetically pleasing compositions. Pre-trained on a curated dataset of high-resolution landscape images and fine-tuned for detailed output, LPGen consistently outperforms existing models in both structural accuracy and stylistic coherence. This work not only advances the field of AI-generated art but also bridges technology with traditional artistic practices, offering valuable insights for future research. The public release of our code, dataset, and model weights will enable broader exploration and application of LPGen in creative domains.